OpenStudy (anonymous):

write an equation in slope intercept form of the line that passes through these points: (3,-2)(6,1) **explain

OpenStudy (anonymous):

no help

OpenStudy (anonymous):

first you need the slope of the line

OpenStudy (anonymous):

points are (3,-2), (6,1) from 3 to 6 is right 3 units, from -2 to 1 is up 3 units right 3, up 3, slope is \(\frac{3}{3}=1\)

OpenStudy (anonymous):

you can also compute via \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{1+2}{6-3}=\frac{3}{3}=1\] if you like

OpenStudy (anonymous):

then write in "point slope form"\[y-y_1=m(x-x_1)\] in your case you have \[y-1=1(x-6)\] or just \[y-1=x-6\] add 1 to solve for \(y\) and get \[y=x-5\]

OpenStudy (anonymous):

@satellite73 why did you do "1+2" if its y2-y1?

OpenStudy (anonymous):

good point \[\frac{1-(-2)}{6-3}\] might have made it clearer i skipped that step

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wouldn't your y-y1 be y-(-2)? @satellite73

OpenStudy (anonymous):

nevermind you did y1 for the second points

OpenStudy (anonymous):

once you have the slope, you can use either point

OpenStudy (anonymous):

right try it with the other point and you will see you get the same equation doesn't matter which one you use